OpenStudy (anonymous):

You took 75 coins to the bank. The coins were a mix of pennies and dimes. If the teller told you there was $4.80 in the pile, how many dimes did you have? 5 15 30 45

OpenStudy (anonymous):

45

OpenStudy (anonymous):

45 * $0.10 = 4.50 the remaining 30 cents is 30 coins so they are pennies

OpenStudy (anonymous):

so is it 45 or 30?

OpenStudy (anonymous):

its 45 dimes

OpenStudy (anonymous):

It would be beneficial to create a system of equations x= penny, y=dime 1.) x+y=75 2.) .01x+.1y=4.80 I preferably dont like decimals so multiply all the terms of the second equation by 100 and get x+10y=480, and then solve for x and get x=480-10y.Then by using substitution you'll get 480-10y+y=75 and combine like terns to get 9y=405 and y=45. So there are 45 dimes.